# Customizing Spider Silk: Generative Models with Mechanical Property Conditioning for Protein Engineering

### v1

@@ 6 Conclusion and Future Work
In parallel, another key focus of future research will be the enhancement of the generative model itself. This will involve integrating significantly larger and more diverse datasets to improve the predictive accuracy and generalizability of the model. By scaling up the data input, the model can better capture the complex relationships between sequence composition and resultant material properties, thereby refining its ability to generate optimized designs. The current methodology, while effective, was specifically designed and optimized for a relatively small dataset. Moving forward, we plan to broaden the scope of this approach by expanding the spidroin-to-mechanical-properties dataset. This expansion will include a wider range of spidroin variants and their corresponding mechanical attributes, enabling a more robust mapping of sequence-to-function correlations and supporting the development of advanced materials with tailored performance characteristics.
Impact Statement
Our research on computational protein design holds transformative potential for developing advanced biomaterials with applications in medicine, sustainability, and biotechnology. By bridging machine learning with protein engineering, we offer a powerful computational framework that can accelerate the design of novel materials with precise mechanical properties. The methodology not only advances scientific understanding but also presents opportunities for addressing critical challenges in sustainable material development and medical innovation, while maintaining a responsible approach to technological advancement.
@@ A.1 Distillation Data Acquisition
A.2 Molecular Weight Calculation
Molecular weight determination is a critical parameter in protein characterization, providing essential insights into protein structure and function. In our research, we employed a standardized approach to molecular weight calculation using established biochemical methods.
The molecular weight of protein sequences was calculated using the standard amino acid molecular weights, accounting for the molecular mass of each amino acid and subtracting the mass of water molecules released during peptide bond formation. Specifically, we utilized the average molecular weights of amino acids as defined by the International Union of Pure and Applied Chemistry (IUPAC) standard [54].
The calculation followed the formula:
Molecular Weight=(Amino Acid Molecular $\text{Molecular Weight}=\sum(\text{Amino Acid Molecular Weight})-(n-1)*18.015$ (5)
Where:
  • $\sum$ represents the sum of individual amino acid molecular weights
  • $n$ represents the number of amino acids in the sequence
  • $18.015$ accounts for the water molecule lost during peptide bond formation
Our computational approach leveraged established bioinformatics libraries to ensure precise and consistent molecular weight calculations across diverse protein sequences.
@@ A.3 Standard Amino Acids with their Background Frequency & KL Divergence for MaSp Repeats
Like human language, protein sequences can be represented as strings of letters, where the protein alphabet consists of 20 standard amino acids (AAs), excluding rare and unconventional ones. Similarly, naturally evolved proteins are composed of modular elements with slight variations, which can be rearranged and assembled hierarchically. In this analogy, common protein motifs and domains—fundamental functional units of proteins—are akin to words, phrases, and sentences in human language [55].

### v2

@@ 6 Conclusion and Future Work
A key direction for future work involves enhancing the generative model by integrating significantly larger and more diverse datasets to improve both predictive accuracy and generalizability. Scaling up the training data will allow the model to better capture the complex, non-linear relationships between sequence composition and resulting mechanical properties, thereby refining its capacity to generate optimized MaSp repeat designs. While the current methodology demonstrates strong performance, it was specifically tailored to a relatively small, high-quality dataset.
To extend this framework, we plan to expand the spidroin-to-mechanical-properties dataset by leveraging a large corpus of unlabelled spidroin sequences available through the Spider Silkome resource (Arakawa et al., 2022). These sequences will be annotated with experimentally or computationally derived mechanical properties in future work, enabling their integration into the training pipeline. This expanded dataset will encompass a broader diversity of spidroin variants and their corresponding mechanical attributes, supporting the development of more robust sequence-to-function mappings and facilitating the design of advanced silk-based materials with tailored performance characteristics.
Software and Data
The implementation of our model inference pipeline is available at: https://github.com/Anon-GitAI/spiderfiber-demo. The repository includes the necessary scripts, configuration files, and usage instructions. All relevant datasets used in this study are provided in the Data folder within the repository.
Acknowledgments
This work was supported by grants from the SciLifeLab & Wallenberg Data Driven Life Science Program (KAW 2020.0239) and by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation: WASP-DDLS 22:035 to AR and HK, KAW 2023.0331 and KAW 2017.0003 to AR, and by the National Bioinformatics Infrastructure Sweden (NBIS) at SciLifeLab. Support was also received by AR from FORMAS (2023-01313), Olle Engkvist stiftelse (233-0334), and the Swedish Research council (2024-02919), and by HK from the Swedish e-Science Research Centre (SeRC).
@@ A.1 Distillation Data Acquisition
A.2 Choice of Physiochemical Attributes for Evaluation
In current study we have used sequence length, molecular weight, instability index, and isoelectric point as physiochemical attributes to evaluate the quality of generated MaSp repeats in various contexts.
Molecular weight and sequence length are widely used as coarse-grained indicators of biological plausibility in protein design. Empirically, naturally occurring major ampullate spidroin (MaSp) proteins fall within a narrow range of lengths and molecular weights, reflecting functional and structural constraints (Yang et al., 2016). For instance, MaSp2 proteins often exhibit high molecular weights due to their repetitive sequences, which are crucial for their mechanical properties (Sino Biological, 2024). Deviations from these ranges can indicate sequences that are unlikely to fold correctly or be expressed efficiently, as proper folding is essential for protein functionality .
The instability index, introduced by (Guruprasad et al., 1990) Guruprasad et al. (1990), predicts the in vivo stability of a protein based on its dipeptide composition. In the context of MaSp repeats, evaluating the instability index helps in predicting the stability of the designed sequences, which is crucial for their functional expression and potential applications.
The isoelectric point is the pH at which a protein carries no net electrical charge. It significantly influences a protein’s solubility, stability, and interaction with other molecules. Understanding the pI of MaSp repeat sequences aids in anticipating their behavior under different pH conditions, which is vital for their purification and application processes (Audain et al., 2016).
In summary, sequence length, molecular weight, instability index and isoelectric point are critical parameters in the evaluation of MaSp repeat sequences, providing insights into their physical constraints, stability and functional suitability in various applications.
@@ A.3 Standard Amino Acids with their Background Frequency & KL Divergence for MaSp Repeats
Like human language, protein sequences can be represented as strings of letters, where the protein alphabet consists of 20 standard amino acids (AAs), excluding rare and unconventional ones. Similarly, naturally evolved proteins are composed of modular elements with slight variations, which can be rearranged and assembled hierarchically. In this analogy, common protein motifs and domains—fundamental functional units of proteins—are akin to words, phrases, and sentences in human language (Ofer et al., 2021).